OpenStudy (anonymous):

Simplify the expression. (5 + 2)[6 – (3 + 2)]

OpenStudy (anonymous):

@zzumfelde009

OpenStudy (texaschic101):

(5 + 2) [ 6 - (3 + 2) ] 7 [ 6 - 3 - 2 ] 7 (1) = your answer

OpenStudy (anonymous):

what my answer?

OpenStudy (texaschic101):

whats 7 times 1 ?

OpenStudy (anonymous):

7

OpenStudy (texaschic101):

thats your answer

OpenStudy (anonymous):

so 7 is my answer

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

oh ok thx

OpenStudy (mathstudent55):

(5 + 2)[6 – (3 + 2)] Start working on the innermost parentheses. = (5 + 2)(6 - 5) Now work on both sets of parentheses since they are at the same level: = (7)(1) Now multiply: = 7